OpenStudy (anonymous):

if log2=0.301 and log3=0.477, then the number of digits in 6^20 is

OpenStudy (campbell_st):

well using log laws 1st deal with the exponent \[\log(x^a) = alog(x) \] apply the above law to the exponent of 20 and the law for multiplication is \[\log(xy) = \log(x) + \log(y) \] where x and y are factors of xy. does this make sense.... ?